On January 1,2020, Ex-B Co. acquired 25% interest in the ordinary shares of KZ Inc. for ₱3,000,000 which reflected book value as of that date. KZ Company reported net income and paid dividends for 2020 and 2021 as follows:

2020 2021

Net income ₱1,000,000 ₱1,500,000

Dividend declared 400,000 700,000

Assume the tax rate is 30% and the inventory is an ordinary asset (i.e., not subjected to final tax).


Case No. 1: On November 20,2020 Ex-B Company sold inventory costing ₱60,000 to KZ Co. for ₱110,000, 60% of which was still unsold on December 31,2020.

1. Investment income in 2020

2. Investment income in 2021

Case No. 2: On November 20,2020 KZ Company sold inventory costing ₱70,000 to Ex-B for ₱120,000, 60% of which was still unsold on December 31,2020.

3. Investment income in 2020

4. Investment income in 2021



1
Expert's answer
2021-06-17T09:32:58-0400

Solution:

Case No. 1:

1.). Investment income in 2020:

Net income 2020 = 1,000,000

Share of net income = 1,000,000 "\times" 30% = 300,000

Adjust for the unsold inventory:

Remaining inventory = 60,000 "\times" 60% = 36,000


Margin = "\frac{110,000 - 60,000}{60,000} = 83.33\%"

Gross profit in ending inventory = 36,000 "\times" 83.33"\%" = 29,999

Unrealized gross profit = 29,999 "\times" 30% = 9,000

Deduct this from the share of net income = 300,000 – 9,000 = 291,000

Investment income in 2020 = 291,000


Investment income in 2021:

Net income 2021 = 1,500,000

Share of net income = 1,500,000 "\times" 30"\%" = 450,000

Add back 2020 unrealized gain now realized = 9,000

Total share of net income = 450,000 + 9,000 = 459,000

Investment income in 2021 = 459,000


Case No. 2:

1.). Investment income in 2020:

Net income 2020 = 1,000,000

Share of net income = 1,000,000 "\times" 30"\%" = 300,000

Adjust for the unsold inventory:

Remaining inventory = 70,000 "\times" 60"\%" = 42,000


Margin = "\frac{120,000 - 70,000}{70,000} = 71.43\%"

Gross profit in ending inventory = 42,000 "\times" 71.43"\%" = 30,001

Unrealized gross profit = 30,001 "\times" 30% = 9,000

Deduct this from the share of net income = 300,000 – 9,000 = 291,000

Investment income in 2020 = 291,000

 

Investment income in 2021:

Net income 2021 = 1,500,000

Share of net income = 1,500,000 "\times" 30"\%" = 450,000

Add back 2020 unrealized gain now realized = 9,000

Total share of net income = 450,000 + 9,000 = 459,000

Investment income in 2021 = 459,000